OpenStudy (gabylovesyou):

Imagine you are teaching a fellow student how to solve: 3 − 5(x + 2) = 3(x + 1) + 6 In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

3 − 5(x + 2) = 3(x + 1) + 6 3- 5x -10 = 3x + 3 +6 -5x -7 = 3x +9 -5x-3x = 9+7 - 8x = 16 8 x = -16 x= ??

OpenStudy (gabylovesyou):

x = -2

OpenStudy (anonymous):

right :)

OpenStudy (anonymous):

now can u see steps and go on?

OpenStudy (gabylovesyou):

for my answer i put that and then i just put x = -2?